At the Summer Olympics, basketball is played on a rectangular court. The width of the land is 1 m more than half of its length. Knowing that the perimeter of the field is 86 m, what are the dimensions of an Olympic basketball court?
Answer by ewatrrr(24785)


 
Hi,
.
width of the land is 1 m more than half of its length (x).
Let x be the Length , width is (x/2) + 1)
P = 2w + 2L
86 = 2(x/2 + 1 ) +2x
86 = x + 2 + 2x
84 = 3x
x = 28m and L = 14+1 = 15m
width is 28m and length is 15m
Checking our Answer
P = 56m + 30m = 86m checks
Wish You the Best in your Studies.